# Active Fine-Tuning of Generalist Policies

### v1

@@ Appendix F Uncertainty ablation
Figure 12: Additional results for AMF-NN in FrankaKitchen with state inputs and different uncertainty quantification techniques. Task allocation during pre-training is reported under each plot.
0510Rounds0.550.600.650.700.75Multi-task success rateMetaworld, uniform
$[4,4,4,4]$
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
$[8,8,0,0]$
0510Rounds0.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
$[0,8,8,0]$
0510Rounds0.50.60.70.8Multi-task success rateMetaworld, skewed
$[0,0,8,8]$
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateMetaworld, skewed
$[16,0,0,0]$
  Last-layer gradient    Last-layer    Dropout
Figure 13: Additional results for AMF-NN in Metaworld with state inputs and different uncertainty quantification techniques. Task allocation during pre-training is reported under each plot.
Appendix G Mitigating forgetting
@@
Several methods aimed at mitigating forgetting can be traced back to rehearsal (Riemer et al., 2019; Chaudhry et al., 2019) and regularization (Kirkpatrick et al., 2017) strategies. While rehearsal approaches are often effective, they also require access to pre-training data, which is unrealistic in our setting. Hence we consider two common regularization technique, namely L2-regularization to the pre-trained weights, and EWC (Kirkpatrick et al., 2017). The latter can be seen as a more nuanced version of the former, which adaptively scales the regularization strength according to the curvature of the loss landscape. Unfortunately, we found both methods to be insufficient in our setting, as reported in Figure 14. When coupled with large regularization weights, their asymptotical performance is significantly limited. When regularization weights are too low, they recover the performance of a naive baseline. Intermediate values were found to interpolate between the two behaviors, without addressing the forgetting issue. While less scalable, we found hard parameter isolation to be the only effective solution amongst the one we tested. As mitigating forgetting in neural network is an orthogonal direction to the main topic of this work, we adopt this solution, and expect that further developments in continual learning will be applicable in our setting.
0510Rounds0.40.50.60.70.8Multi-task success rateKitchen, uniform
0510Rounds0.20.30.40.50.60.70.8Multi-task success rateKitchen, skewed
0510Rounds0.40.50.60.7Multi-task success rateMetaworld, uniform
0510Rounds0.20.40.6Multi-task success rateMetaworld, skewed
  Parameter isolation    EWC    L2 regularization    Baseline
@@
Figure 14: Performance of uniform task sampling with several techniques to mitigate forgetting. Darker shades represent stronger regularization coefficients. L2 and EWC regularization are not effective in this setting.
Appendix H Criterion vs returns
@@
Appendix I Implementation Details
In order to ease reproducibility, we open-source our codebase on the project’s anonymous website.444sites.google.com/active-multitask-finetuning Furthermore, we describe several implementation details in the following sections.
I.1 Metrics
@@ I.2 GP settings
I.3 Neural network settings
I.3.1 Environments
We evaluate AMF-NN across three environments, namely FrankaKitchen, Metaworld and WidowX. For the first two, demonstrations are $50$ steps, while for the latter they involve $100$ steps. In FrankaKitchen, demonstrations are provided by Kumar et al. (2024), and collected by trained SAC agents. In Metaworld, demostrations are instead collected by the scripted policies provided (Yu et al., 2020). Finally, in WidowX successful trajectories are collected by Octo-small (Octo Model Team et al., 2024) itself and filtered according to success labels, in an instance of self-supervised distillation. Furthermore, in the case of WidowX, the initial position of the object is not randomized, as we found this to result in very inconsistent performance for the data collection policy. In the first two environments, $25$ attempts for each task are evaluated, while the evaluation for WidowX involves $50$ attempts.
I.3.2 Behavior cloning
The policy is parameterized via a deterministic MLP with $2$ layers and $256$ units per layer, with layer normalization (Ba, 2016). Task conditioning are image embedding extracted by R3M (Nair et al., 2022a). Policies are pre-trained for $500$ epochs with batch size of $128$, learning rate of $41.e-$ 4 using the Adam optimizer (Kingma, 2014).
I.3.3 Importance sampling weights
@@
I.3.4 AMF
Each fine-tuning round involves $3000$ gradient steps, each with a batch size of $128$. We warm-start each algorithm by collecting the first $|\mathcal{C}|$ demonstrations uniformly, as mentioned in Section 4.2. In the case of loss-gradient embeddings, we found it to be beneficial to use a separate copy of the policy for task selection, which is not trained on these initial trajectories (which thus can be seen as a small “validation” set). As these demonstrations are not selected according to the criterion, to avoid unwanted updates of pre-training weights, they are only added to the training set once the algorithm selects a task belonging to the same partition of the task space, as defined by parameter isolation.
Appendix J Useful inequalities
Lemma 8.

### v3

@@ Appendix G Uncertainty ablation
Figure 11: Additional results for AMF-NN in Metaworld with state inputs and different uncertainty quantification techniques.
Section 5.3 evaluates alternative uncertainty quantification schemes in FrankaKitchen for two representative pre-training distributions. This Section extends these results to include results for Metaworld (see Figure 11). Results are consistent with those so far reported, suggesting that loss gradient embeddings are an important component for the empirical performance of AMF-NN.
Appendix H Mitigating forgetting
@@
01020Rounds0.20.40.60.8Multi-task success rateKitchen, mismatch
01020Rounds0.600.650.700.750.800.85Multi-task success rateKitchen, no mismatch
01020Rounds0.20.40.60.8Multi-task success rateMetaworld, mismatch
01020Rounds0.30.40.50.6Multi-task success rateMetaworld, no mismatch
  Adaptive Prior    EWC    L2 regularization    Baseline
@@
Figure 12: Performance of AMF-NN with several techniques to mitigate forgetting. Darker shades represent stronger regularization coefficients. L2, EWC regularization are not effective in this setting.
Several methods aimed at mitigating forgetting can be traced back to rehearsal (Riemer et al., 2019; Chaudhry et al., 2019) and regularization (Kirkpatrick et al., 2017) strategies. While rehearsal approaches are often effective, they also require access to pre-training data, which is unrealistic in our setting. Hence we consider two common regularization technique, namely L2-regularization to the pre-trained weights, and EWC (Kirkpatrick et al., 2017). The latter can be seen as a more nuanced version of the former, which adaptively scales the regularization strength according to the curvature of the loss landscape. Unfortunately, we find these methods to be insufficient in our setting, as reported in Figure 12. When coupled with large regularization weights, the asymptotical performance of L2-regularization and EWC is significantly limited. When regularization weights are too low, they recover the performance of a naive baseline. Intermediate values were found to interpolate between the two behaviors, without addressing the forgetting issue. As a result, the policy quickly loses information on its pre-training tasks, thus rendering adaptive data selection strategies ineffective.
This motivates our adoption of a novel technique to alleviate forgetting, which we refer to as Adaptive Prior. Inspired by previous works in behavioral priors (Bagatella et al., 2022) and offline RL (Kumar et al., 2020), we linearly combine the policy’s output with that of a prior (in practice, a frozen copy $\bm{\pi}_{p}$ of the pre-trained policy). When an action needs to be sampled, we instead output a linear combination of actions sampled from the fine-tuned policy $\bm{\pi}$ and the prior $\bm{\pi}_{p}$: for a given state-task pair $(s,c)\in\mathcal{S\times C}$, the action selected is $a=\alpha(c)\hat{a}+(1-\alpha(c))\bar{a}$, where $\hat{a}\sim\bm{\pi}(\cdot|s,c)$ and $\bar{a}\sim\bm{\pi}_{p}(\cdot|s,c)$. The weight $\alpha(c)\in[0,1]$ is task-dependent and learnable: ideally, it would be high when the fine-tuned policy is more suitable for the task, and low when the prior is instead preferable. In practice, it may be a per-task learnable parameter for finite task spaces, or the output of a parametrized function (e.g., a neural network) in continuous task spaces. While this formulation departs from existing ones (Bagatella et al., 2022), in which the action is instead sampled from a mixture of policy and prior, it enables a simple update rule for $\alpha$, which remains applicable when the policy class does not allow straightforward likelihood evaluation (e.g., Diffusion policies). Under the assumption that both $\bm{\pi}$ and $\bm{\pi}_{p}$ are isotropic Gaussians, the BC loss computed with respect to the linear combinations of actions simplifies to a simple mean squared error:
$\mathcal{L}^{\alpha}_{\text{BC}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=0}^{H-1}\|a_{t}^{i}-(\alpha(c_{i})\hat{a}+(1-\alpha(c_{i}))\bar{a})\|_{2},$ (93)
where $\hat{a}\sim\bm{\pi}(\cdot|s_{t}^{i},c_{i}),\bar{a}\sim\bm{\pi}_{p}(\cdot|s_{t}^{i},c_{i})$ and $\hat{\tau}_{1:N}=(s_{0}^{i},a_{0}^{i},\dots,s_{H-1}^{i},a_{H-1}^{i})_{i=1}^{N}$ is the dataset of $N$ task-conditioned, $H$-length trajectories with task labels $c_{1:N}$. As this loss function is differentiable with respect to $\alpha(c)$, the weight can be easily trained through gradient descent. Intuitively, this gradient pushes weights up for tasks in which the fine-tuned policy is more accurate than the pre-trained one. As $\bm{\pi}$ is fine-tuned on the dataset used for estimating $\mathcal{L}_{\text{BC}}^{\alpha}$, its performance can be easily overestimated. Thus, we propose to introduce a conservative penalty term, which encourages the weight to be updated only if the fine-tuned policy $\bm{\pi}$ significantly outperforms the prior $\bm{\pi}_{p}$: $\mathcal{L}^{\alpha}=\mathcal{L}^{\alpha}_{\text{BC}}+\beta\frac{1}{N}\sum_{i=1}^{N}\alpha(c_{i}).$ We remark that, in our empirical evaluation, the Gaussian assumption is often violated; nevertheless, we found this update rule to be empirically effective across models and tasks. To illustrate, we visualize trajectories of mixing weights $\alpha(c)$ during fine-tuning in Figure 13. We consider a mismatched setting, in which the pre-training distribution largely demonstrates the first two tasks in Kitchen, and does not cover the remaining tree. During early training, AMF-NN mostly samples demonstrations for the last three tasks. Thus, $\bm{\pi}$ quickly improves, and the respective mixing coefficients converge to 1. On the other hand, $\bm{\pi}_{p}$ outperforms $\bm{\pi}$ on the first two tasks for the majority of training, $\alpha$ remains low and evaluation performance does not drop as actions are largely sampled from $\bm{\pi}_{p}$.
[…]
Figure 13: Success rates (blue) and mixing coefficients $\alpha(c)$ (grey) over training in Kitchen with skewed pre-training. $\alpha$ grows quickly for novel tasks, and more conservatively for those that were demonstrated during pre-training.
Appendix I Does AMF rebalance demonstration counts?
In discrete task spaces, counting the number of demonstrations for each task is possible. In this case, a naive data selection strategy would simply request demonstrations for tasks that have been demonstrated the least in the past. If all tasks require a similar amount of demonstrations, this would empirically perform very well. In our setting, however, data selection algorithms do not have knowledge of pre-training data. For this reason, a count could only be kept with respect to the fine-tuning demonstrations: actively balancing this count would lead to a near-uniform task selection, and recover the performance of uniform sampling in expectation.
01020Rounds0.20.40.60.8Multi-task success rateKitchen, mismatch
01020Rounds0.600.650.700.750.800.85Multi-task success rateKitchen, no mismatch
01020Rounds0.20.40.60.8Multi-task success rateMetaworld, mismatch
01020Rounds0.30.40.50.6Multi-task success rateMetaworld, no mismatch
  AMF-NN    Rebalancing    Uniform, with adaptive prior    Uniform
Figure 14: Extended results from Figure 4, including a privileged “rebalancing” baseline.
Nevertheless, we implement this “rebalancing” criterion as a privileged baseline, which assumes access to the pre-training task distribution. We evaluate it in the standard settings for AMF-NN from Figure 4. In Figure 14, we observe that AMF-NN is able to match the performance of this privileged baseline, despite having no knowledge of the pretraining distribution.
This implies that AMF can infer information on the pre-training phase through estimation of the policy’s uncertainty, and is capable of automatically recovering a “rebalancing” strategy. Moreover, AMF-NN considers the reduction in entropy across several tasks: hence, it can in contrast focus on tasks that are harder to learn or that could, in principle, lead to learning progress on other tasks. Further empirical evidence for these behaviors is shown in Appendix J.
Appendix J Single-task performance
This Section presents a detailed look at the data selection strategies induced by AMF-NN. For this purpose, we consider the main experiments in Kitchen and Metaworld outlined in Figure 4, and plot single-task success rates, as well as the amount of demonstrations collected over time. For reference, asymptotic single-task success rates on Metaworld and Kitchen converge between 70% and 100%, depending on the task.
In the case of skewed pre-training (Fig. 16 and 18), we observe that AMF samples tasks that were not present in the pretraining dataset more often, without having access to any direct information on the pre-training distribution. Moreover, even if multiple tasks have the same frequency in the pre-training distribution, AMF will prefer the ones that induce a larger reduction in posterior uncertainty: for instance, in Metaworld, AMF selects Close Faucet more often than Open Faucet, despite the fact that both were similarly demonstrated during pre-training. We remark that these task selection strategies arise naturally from our information-based criterion in Equation 1, without any direct information on the pre-training distribution, nor any explicit policy evaluation.
  AMF-NN ($\sigma=1e^{-3}$)   Uniform, with adaptive prior    # of pre-training demonstrations
Figure 15: Single-task curves for skewed pre-training in Kitchen. Dashed lines represent demonstrations counts, with grey lines displaying the (inaccessible) count of pre-training demonstrations.
  AMF-NN ($\sigma=1e^{-3}$)   Uniform, with adaptive prior    # of pre-training demonstrations
Figure 16: Single-task curves for uniform pre-training in Kitchen. Dashed lines represent demonstrations counts, with grey lines displaying the (inaccessible) count of pre-training demonstrations.
  AMF-NN ($\sigma=1e^{-3}$)   Uniform, with adaptive prior    # of pre-training demonstrations
Figure 17: Single-task curves for skewed pre-training in Metaworld. Dashed lines represent demonstrations counts, with grey lines displaying the (inaccessible) count of pre-training demonstrations.
  AMF-NN ($\sigma=1e^{-3}$)   Uniform, with adaptive prior    # of pre-training demonstrations
Figure 18: Single-task curves for uniform pre-training in Metaworld. Dashed lines represent demonstrations counts, with grey lines displaying the (inaccessible) count of pre-training demonstrations.
Appendix K Analysis of Importance Weights
Importance weights (as introduced in Equation 3) allow estimating the expert’s occupancy for arbitrary tasks. Naturally, the quality of importance weights depends on many factors, including the dimensionality of the trajectory space, and the density with which available data covers it. In this section, we report a qualitative evaluation of importance weights for both AMF-GP and AMF-NN (Figures 19 and 20, respectively). In both cases, we find that informative weights can be retrieved eventually, given the proper amount of clipping (as described in Appendix M). While in early rounds of the algorithm, weights can be inaccurate, leading to a poor estimate of the objective, we observe that the quality of importance sampling weights improves within a handful of rounds.
Figure 19: Analysis of importance sampling weights for AMF-GP. We consider the skewed pre-training setting from Figure 2, and compute importance weights after 1, 5 and 9 rounds. We sample four tasks $c_{0:3}$, represented by vertical dashed lines of different colors. For each task $c_{i}$, we collect a demonstration $\tau_{i}$ and sweep over $c^{\prime}\in\mathcal{C}$ on the x-axis; we plot $w(\tau_{i},c^{\prime})$ with solid lines. We observe that importance weights are uninformative in early parts of training, but converge to more accurate values within a few rounds.
Figure 20: Analysis of importance sampling weights for AMF-NN. We consider the skewed pre-training setting from Figure 4, and compute importance weights after 1, 10 and 19 rounds. We visualize weights for both Kitchen (top) and Metaworld (bottom). As the task set is discrete, we consider all tasks $(c_{i}\in\mathcal{C})$, and collect one demonstration $\tau_{i}$ for each. The entry of each colormap at row $i$ and column $j$ represents $w(\tau_{j},c_{i})$. Again, we observe that at the beginning of training importance weights can be inaccurate, particularly for tasks $c_{i}$ that have not been sufficiently demonstrated. However, as more data is collected and the policy specializes to each task, the weights converge.
Appendix L Criterion vs returns
@@ Appendix I Does AMF rebalance demonstration counts?
Appendix M Implementation Details
In order to ease reproducibility, we open-source our codebase on the project’s repo.444github.com/marbaga/amf Furthermore, we describe several implementation details in the following sections.
M.1 Metrics
@@
M.3 Neural network settings
M.3.1 Environments
We evaluate AMF-NN across four environment suites, namely FrankaKitchen, Metaworld, Robomimic and WidowX. For the first two, demonstrations are $\approx 50$ steps, while while for the latter two they involve up to $700$ steps, and $100$ steps respectively. In FrankaKitchen, demonstrations are provided by Kumar et al. (2024), and collected by trained SAC agents. In Metaworld, demostrations are instead collected by the scripted policies provided (Yu et al., 2020). In Robomimic, demonstrations are provided by proeficient human demonstrators (Mandlekar et al., 2021). Finally, in WidowX successful trajectories are collected by Octo-small (Octo Model Team et al., 2024) itself and filtered according to success labels, in an instance of self-supervised distillation. Furthermore, in the case of WidowX, the initial position of the object is not randomized, as we found this to result in very inconsistent performance for the data collection policy. In the first two suites, $50$ attempts for each task are evaluated, while evaluation on Robomimic and WidowX involves $25$ attempts.
M.3.2 Behavior cloning
The MLP policy has with $2$ layers and $256$ units per layer, with layer normalization (Ba, 2016). The Diffusion policy shares architecture and hyperparameters with the original implementation (Chi et al., 2023). Task conditioning are word embedding extracted by a sentence transformer (all-MiniLM-L6-v2) (HuggingFace, 2026). Policies are pre-trained for $200$ epochs with batch size of $256$, learning rate of $10^{-4}$ using the AdamW optimizer(Loshchilov & Hutter, 2019).
M.3.3 Importance sampling weights
@@
M.3.4 AMF
Each fine-tuning round involves $3000$ gradient steps, each with a batch size of $256$. We warm-start each algorithm by collecting the first $|\mathcal{C}|$ demonstrations uniformly, as mentioned in Section 4.2. In the case of loss-gradient embeddings, we found it to be beneficial to use a separate copy of the policy for task selection, which is not trained on these initial trajectories (which thus can be seen as a small “validation” set).
M.3.5 Adaptive Prior
We found relatively high learning rates for the mixing weight $\alpha$ to work well: all experiments use a learning rate of $1.0$. The conservative coefficient $\beta$ was tuned for each suite, and is set to $0.01$ for all but Kitchen, in which it takes the value of $0.001$.
M.4 Runtime
Each experimental run for AMF-NN takes at most 5 hours with GPU acceleration. In this case, data selection itself requires less than 1 second per round, and can be significantly sped up by reducing the sampling budget. AMF-GP experiments can be reproduced within 10 minutes on CPU.
Appendix N Useful inequalities
Lemma N.1.